OpenStudy (anonymous):

Find the equation of the line that contains the given points. (Let y be the dependent variable and let x be the independent variable.)p1(4,5) p2(3,-4) could someone help me please?

OpenStudy (anonymous):

Do you recall the formula for the slope of a line?

OpenStudy (anonymous):

given two points?

OpenStudy (amistre64):

find the slope by subtracting one set of points from the other olike this: (4, 5) -(3,-4) (4, 5) (-3,4) ------- (1,9) slope = y/x = 9/1 = 9

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (amistre64):

( x , y ) -(x1,y1) -------- yes....

OpenStudy (anonymous):

ok thanks!

OpenStudy (amistre64):

i find it easier to stack them; and subtract the bottom from the top.... then slope = m = y/x

OpenStudy (anonymous):

you know I haven't thought of that . It looked too easy!

OpenStudy (amistre64):

so we have an equation now that looks like: y = 9x + b ; plug in one of the points to solve for b

OpenStudy (amistre64):

5 = 9(4) + b 5 = 36 + b 5 -36 = b -31 = b .................... y = 9x -31

OpenStudy (amistre64):

another format is to put all the variables to one sidelike this: -9x +y = -31 ; but some textbooks hate to have a (-) out front so they do this: 9x -y = 31

OpenStudy (anonymous):

ok the first way looks easier!

OpenStudy (amistre64):

:) y = 9x-31 is the "slope,intercept" form. 9x-y=31 is the "standard" form